A coupon bond paying semiannual interest is reported as having an ask price of 105% of its $1,000 par value. If the last interest payment was made 60 days ago and the coupone rate is 496, what is the invoice price of the bond? Assume there are 182 days between coupon payments. Round your answer to two decimal places, and enter it without a dollar sign. Your Answer: Answer Question 9 (1 point) Consider a bond that matures in 7 years. The bond has a par value of $1,000, 8% semiannual coupon. what is the bonds price today if the interest rate is 7.0%? Round your answer to two decimal places. Your Answer: Answer
Answer #1

8). Semi-annual coupon = 1,000 x .04 x .5 = $20.

One month of accrued interest is $20 x (60/182) = $6.593

at a price of 105% the invoice price is 1,050 + 6.593 = $1,056.59

9). To find the bond price, we need to put the following values in the financial calculator:

INPUT 7x2=14 7/2=3.5 (8%/2)x1,000=40 1,000
TVM N I/Y PV PMT FV
OUTPUT -1,054.60

Thus, Current Bond's Price = $1,054.60